Moving    26   29.9   39.5   25.1   31.6   36.2   25.1   31 35.6   40.

Statics   27.9 29.1 38 27 30.3 34.5 27.8 29.6 33.1 35.5

Continuing the use of the lorry data above,

(a)   Compute the mean and standard deviation for weight in motion.

(b)   Compute the mean and standard deviation for static weight.

(c)   Compute the correlation coefficient (Note: You have already done this in Exercise 06.TR.A. You need only copy it.)

(d)   Use the formulas on pages 199 – 201 of the text to find the slope of the regression line.

(e)   Use these formulas to find the y-intercept of the regression line. (f)   Use the above two results to write the regression equation?

(g)   Check to see that this agrees with the result of part (a) of Exercise 07.BU.A within rounding.
Answer #1

Find X Y . X and Y ? as it was done in the table below. Y 27.9 Y.Y 778.41 846.81 1444 29.1 729 X-Y 725 4 870.09 1501 677.7 95

Find the sum of every column: ΣΧ = 320, ΣΥ = 312.8, ΣΧ.Υ = 10194.31, Σx? - 10527.64 Use the following equations to find a and

| (a) Mean = x - 320 = [32] n- 1 X 1. (6) Mean = sy = 312-8 = 31.28] so - 121.036 - 3-754] 10 10-) (c) from the above calcula